### 10 Mar 2026 v1

# MiniAppBench: Evaluating the Shift from Text to Interactive HTML Responses in LLM-Powered Assistants Zuhao Zhang1,2*, Chengyue Yu1*, Yuante Li3, Chenyi Zhuang1†, Linjian Mo1, Shuai Li2

Abstract
With the rapid advancement of Large Language Models (LLMs) in code generation, human-AI interaction is evolving from static text responses to dynamic, interactive HTML-based applications, which we term MiniApps. These applications require models to not only render visual interfaces but also construct customized interaction logic that adheres to real-world principles. However, existing benchmarks primarily focus on algorithmic correctness or static layout reconstruction, failing to capture the capabilities required for this new paradigm. To address this gap, we introduce MiniAppBench, the first comprehensive benchmark designed to evaluate principle-driven, interactive application generation. Sourced from a real-world application with 10M+ generations, MiniAppBench distills 500 tasks across six domains (e.g., Games, Science, and Tools). Furthermore, to tackle the challenge of evaluating open-ended interactions where no single ground truth exists, we propose MiniAppEval, an agentic evaluation framework. Leveraging browser automation, it performs human-like exploratory testing to systematically assess applications across three dimensions: Intention, Static, and Dynamic. Our experiments reveal that current LLMs still face significant challenges in generating high-quality MiniApps, while MiniAppEval demonstrates high alignment with human judgment, establishing a reliable standard for future research. Our code is available in github.com/MiniAppBench.
[…]
MiniAppBench: Evaluating the Shift from Text to
Interactive HTML Responses in LLM-Powered Assistants
Zuhao Zhang1,2*, Chengyue Yu1*, Yuante Li3, Chenyi Zhuang1†, Linjian Mo1, Shuai Li2
1Inclusion AI, Ant Group 2Shanghai Jiao Tong University 3Carnegie Mellon University
1 Introduction
With the rapid advancement of Large Language Models (LLMs) in code generation (Novikov et al., 2025; Li et al., 2025c; Xia et al., 2025), models are evolving to Autonomous Architects capable of constructing complete software solutions. In this emerging landscape, code transcends its role as a mere intermediate symbolic representation; it becomes a direct executable medium through which a model’s internal knowledge is externalized into dynamic, user-facing artifacts.
This transformation facilitates a paradigm shift in human-LLM interaction (as illustrated in Figure 1), moving from static text-only responses to rich, code-based engagements. Users now expect LLMs to produce interactive visualizations or functional applications that embody real-world logic. Consequently, to ensure these interactions feel natural and seamless, the model must actively capture and construct implicit assumptions or principles, such as “an object in free fall follows Newton’s laws" or “a week has seven days", which, while often taken for granted in human communication, are essential for valid execution. Real-world cases are shown in Figure 2.
We argue that the web provides a particularly effective substrate for realizing such interactions. In this context, HTML represents world states and structural relationships, CSS determines perceptual salience, and JavaScript encodes causal dependencies, temporal evolution, and interaction logic—together forming an executable world model. Moreover, its interactivity adds an additional layer of depth to this interaction.
From this perspective, we posit that rendered HTML responses will emerge as a new form of human–LLM interaction, which we term MiniApps. Unlike traditional web pages, which primarily focus on static content display or predefined CRUD (Create, Read, Update, and Delete) workflows, MiniApps are characterized by two core properties: ❶ Fidelity to Real-World Principles, where the model must capture and construct the implicit principles embedded in the user’s query; and ❷ Customized Interaction, where application structure and behavior are dynamically synthesized to match user intent, rather than being instantiated from fixed templates.
However, current benchmarks remain tethered to the static past, failing to capture this shift. Traditional code benchmarks like MBPP (Austin et al., 2021) and HumanEval (Chen, 2021) focus on algorithmic syntax, treating code as abstract logic divorced from execution context. Conversely, web generation benchmarks (Sun et al., 2025; Lu et al., 2025; Xu et al., 2025) prioritize visual fidelity or static layout reconstruction. This creates a critical blind spot: existing metrics are unable to verify whether LLMs truly capture and construct the underlying real-world principles implied by user queries.
In practice, achieving these properties is non-trivial. As shown in Figure 2, an artifact may be syntactically valid and successfully executable, but still fail to support high-fidelity, non-fragmented interaction aligned with real user reasoning. To bridge this gap, we introduce MiniAppBench, the first benchmark designed specifically to evaluate the ability of LLMs to generate MiniApps. The comparison with other benchmarks is provided in the Appendix A. MiniAppBench is constructed through a rigorous multi-stage pipeline that distills tens of millions of real-world user queries into a balanced set of principle-driven, interaction-intensive tasks.
Evaluating MiniApps also poses a unique challenge due to the inherently open-ended nature of application generation. Given that multiple implementations with different structures, interaction patterns, and design choices may all validly satisfy the same user intent, there is often no single canonical “ground truth” code solution.
To address this challenge, we propose a novel Agentic Evaluation Framework , MiniAppEval. Instead of relying on rigid assertions or template-based matching, MiniAppEval leverages Playwright (Microsoft, 2026) to perform human-like exploratory testing by simulating interactions such as clicking, dragging, and observing runtime behavior. It dynamically verifies the generated application along three complementary dimensions: Intention, Static, and Dynamic. Together, these dimensions assess whether the application fulfills the user’s intent, exhibits a coherent static implementation, and demonstrates interactive behavior that adheres to implicit real-world constraints and interaction expectations.
Our main contributions are summarized as follows:
We rethink the future of human-LLM interaction and argue that rendered HTML responses constitute a new interaction paradigm in the form of MiniApps.
We propose MiniAppBench, the first benchmark dedicated to evaluating principle-driven, interactive application generation. Derived from real-world user demands, it comprises 500 rigorous tasks that challenge LLMs to align executable code with implicit user reasoning.
We introduce MiniAppEval, a novel agentic framework that integrates static inspection with human-like dynamic exploration to holistically assess application fidelity across Intention, Static, and Dynamic.
Experiments reveal that current LLMs still struggle to reliably construct MiniApps, while MiniAppEval achieves high consistency with human judgment, enabling more faithful assessment of next-generation interactive systems.
2 Related Work
2.1 Code Generation and World Reasoning
Existing code generation benchmarks (Paul et al., 2024; Jiang et al., 2024) have largely focused on assessing functional correctness within the domains of algorithmic logic, software engineering, and data science. Early benchmarks such as HumanEval (Chen, 2021) and MBPP (Austin et al., 2021) assess function-level algorithmic reasoning, while more recent efforts like SWE-bench (Jimenez et al., 2023) and MLE-bench (Chan et al., 2024) extend evaluation to repository-scale software maintenance and engineering workflows. Despite this progression in scale and realism, these benchmarks largely treat code as an abstract symbolic artifact whose quality is determined by test passing or task completion. Interaction and user-facing behavior are either absent or tightly constrained by fixed assertions. As a result, they do not capture whether models can use code as an interactive medium to externalize knowledge, reason about real-world principles, or support customized human-LLM interaction—capabilities that are central to MiniApps.
Conversely, a parallel line of research evaluates LLMs on their understanding of real-world principles. Benchmarks such as PIQA (Bisk et al., 2020) and GSM8K Cobbe et al. (2021) assess this capability through passive textual inference, asking models to predict outcomes based on described scenarios. In the domain of embodied AI, frameworks like AlfWorld Shridhar et al. (2020) and Voyager Wang et al. (2023) test agents’ ability to act within predefined, immutable environments. While these benchmarks explicitly evaluate models’ understanding of explicit real-world principles within constrained scenarios, they do not assess the ability of models to capture and integrate implicit principles and express them through executable artifacts.
2.2 Web Development
Early work on web generation (Li et al., 2025b; Ning et al., 2025) mainly focused on visual-to-code translation and static layout reconstruction. Pioneering works like Pix2Code (Beltramelli, 2018) and Web2Code (Yun et al., 2024) treated web generation as an image captioning or translation task, focusing on pixel-level fidelity and structural alignment with reference designs. Similarly, benchmarks like FullFront (Sun et al., 2025) emphasize the visual consistency of the generated frontend. Sketch2Code (Li et al., 2025b) further extended this to hand-drawn sketches. These approaches largely focus on visual appearance, with limited attention to the dynamic logic and state transitions that characterize modern interactive applications. More recent benchmarks have advanced towards Engineering-level Web Development, addressing multi-step or multi-file generation. Frameworks such as WebGenBench (Lu et al., 2025) and WebBench (Xu et al., 2025) evaluate the ability to construct complex file structures for traditional applications like e-commerce sites or forums. However, despite increased structural complexity, these tasks remain centered on information presentation and standard CRUD workflows, often relying on templates and established patterns, with limited need for reasoning about custom interaction rules.
2.3 Evaluation Methodologies
Traditional web evaluation paradigms typically rely on static code analysis, visual similarity metrics (e.g., screenshot comparison), or predefined interaction scripts. Approaches like Pix2Code (Beltramelli, 2018) and Web2Code (Yun et al., 2024) adopt snapshot-based evaluation, which captures layout fidelity but overlooks the interaction process. ArtifactsBench (Zhang et al., 2025), on the other hand, analyzes the interaction process through multiple screenshots. Similarly, methods relying on fixed click-scripts, such as WebBench (Xu et al., 2025), FullFront (Sun et al., 2025), cover only narrow, pre-determined paths. In contrast, modern interactive applications feature rich interactivity and effectively unbounded state spaces. Fixed scripts cannot adapt to diverse valid behaviors or open-ended interaction trajectories implemented by a model. Consequently, static or scripted methods are ill-equipped to evaluate whether a generated application truly functions as a consistent dynamic system.
While recent works have introduced agent-based evaluators (Wang et al., 2024; Gao et al., 2024) to address interactivity, they predominantly rely on comparative analysis. Systems like WebDevJudge (Li et al., 2025a) and FronTalk (Wu et al., 2025) evaluate quality by measuring deviation from a reference implementation (ground truth) or by performing pairwise preference rankings (A/B testing). Such reference-dependent evaluation is ill-suited for MiniApps, where customized and open-ended generation admits multiple equally valid realizations.
3 MiniAppBench
3.1 Overview
We present MiniAppBench, a benchmark comprising 500 tasks designed to evaluate LLMs on their ability to develop MiniApps as a new form of human-LLM interaction. Moving beyond static layouts or standard CRUD operations found in prior work Xu et al. (2025); Zhang et al. (2025), our benchmark focuses on adherence to real-world principles and customized interaction. The dataset is distilled from tens of millions of real user queries collected from a large-scale production platform. Through a multi-stage filtration process involving model-based difficulty assessment and manual verification (detailed in Appendix B), we selected 500 high-value queries that span six diverse domains (see Figure 3(e)). Critically, these tasks require models not only to generate syntactically valid code, but also to construct interactive behaviors that align with user intent by correctly capturing and operationalizing implicit real-world principles, thereby enabling coherent, natural, and non-fragmented user interactions. The overview of MiniAppBench is provided in Figure 3.
3.2 Data Representation
To facilitate structured evaluation and fine-grained analysis, we organize the dataset into a canonical tuple representation. Formally, the dataset is defined as , where each entry is encapsulated as:
Here, the components are defined as follows (the data format is described in Appendix B.3):
represents the natural-language query sourced from real users, serving as the input for the model.
denotes the two-level taxonomy, where is the coarse-grained domain (e.g., Science, Games) and is the specific subclass, enabling domain-specific performance breakdown.
is the structured evaluation reference. Unlike traditional benchmarks that rely on fixed test cases, specifies verifiable constraints across Intention, Static, and Dynamic dimensions to guide the agentic evaluator.
labels the task difficulty, derived from the pass rates of baseline models.
This structured representation supports the open-ended nature of MiniApps: the evaluation reference functions as a flexible inspection guide rather than a rigid template, validating any generated artifact that functionally satisfies the user intent .
3.3 Evaluation Dimension
We design three dimensions to assess the quality of MiniApps, comprehensively verify whether the generated application adheres to the real-world principles and interaction expectations specified by the user.
Intention Dimension.
This score measures whether the MiniApp correctly interprets and fulfills the high-level user goal specified in . For example, if the query requests a physics simulation of pendulum motion, the evaluator checks whether the core dynamics (periodicity, energy conservation) are meaningfully represented.
Static Dimension.
This score evaluates structural and syntactic correctness without execution. It verifies the presence of required elements, proper code organization, and adherence to accessibility standards. For instance, a weather dashboard should include clearly labeled temperature, humidity, and location fields, despite interaction.
Dynamic Dimension.
This score evaluates the MiniApp’s runtime behavior through multi-step interaction trajectories. It evaluates two critical aspects: (1) Sequential Logic and Planning: The evaluator executes complex chains of actions (e.g., add a new task mark as complete verify removal from the active list) to verify that state transitions remain consistent and reversible, faithfully reflecting causal dependencies in the real world. (2) Robustness and Boundary Handling: MiniAppEval is tested against adversarial or edge-case inputs (e.g., submitting an empty string as a task name or inputting invalid dates in a scheduler) to ensure the application handles exceptions gracefully without crashing or violating real-world principles.
3.4 Dataset Construction Pipeline
➠Stage 1: Identifying Principle-Driven Interactive Queries.
The first stage tackles a key challenge: not all real user queries are suitable for evaluating customized interaction or the construction of real-world principles. Many queries are purely informational, underspecified, or trivially solvable without meaningful interaction logic.
We began with an initial pool of tens of millions of real user queries, from which we sampled a subset and removed invalid entries (e.g., incoherent text, multi-turn follow-ups), resulting in 3,234 candidates. We then used a LLM-based categorization approach to group queries by their underlying themes and suitability for interactive tasks. Human experts further refined these categories into 6 coarse-grained domains and 25 fine-grained subclasses, ensuring semantic consistency and balanced coverage across knowledge areas (details in Appendix B.1). To ensure data quality, we applied a hybrid quality filtering strategy. First, an LLM-driven filter removed queries that were vague, static, or lacking in interactive potential. Second, a manual verification step confirmed that the underlying principles and interactive logic of each task could be explicitly materialized through HTML (the full pipeline is provided in Appendix B.2.2). This rigorous verification ensures that every task in the dataset is suitable for testing the core aspects of the benchmark.
This stage resulted in 1,123 high-quality seed queries, forming the foundation of the benchmark. These queries are rich in real-world principles and support meaningful evaluation of customized interactions and principle-based generation.
➠Stage 2: Expanding Coverage While Preserving Core Intent.
While the filtered seed queries are high quality, they alone do not provide sufficient coverage of interaction patterns or domain diversity. The second stage therefore focuses on expanding task diversity without diluting the underlying principles. We employ the seed queries as anchors in an LLM-driven evolutionary augmentation process to synthesize variants. These variants explore diverse scenarios, parameter configurations, and interaction structures while strictly maintaining the original intent. Both seed and generated queries then undergo a standardization step, in which they are rewritten to be self-contained, explicit, and engineering-feasible.
This step is critical, as it ensures the benchmark evaluates application construction ability rather than ambiguity resolution or prompt interpretation. After augmentation and standardization, the query set expands to 1,974 candidates.
➠Stage 3: Anchoring Tasks with Verifiable Evaluation References.
We sampled 200 queries from Stage 2 and asked different models to generate MiniApps for manual assessment. During this process, we identified both cross-domain issues and domain-specific pitfalls. To enhance the evaluation capability of MiniAppEval, we construct evaluation references via a human-guided generation strategy. Specifically, human experts write (i) a set of general guidelines and (ii) domain-specific instructions to guide an LLM in generating these references.
Given the query , its domain , and the guidelines , the LLM maps key evaluation points onto three dimensions aligned with our evaluation dimension and produces a query-specific reference:
These references assist the evaluator but are not used as the final decision criterion. We further asked domain experts to audit the generated reference. Their review suggests that the reference effectively surfaces implicit underlying principles that the MiniApps generation model might otherwise overlook (Figure 2). Importantly, the references are not manually refined, ensuring scalability, generalizability, and full reproducibility.
➠Stage 4: Balancing Difficulty and Domain Coverage.
The final stage constructs a balanced, challenging, and statistically meaningful evaluation benchmark. Tasks are assessed along the Intention, Static, and Dynamic dimensions and categorized into Easy, Medium, or Hard levels. To ensure diversity and fairness, we perform stratified sampling, selecting 500 tasks from a combination of domains and difficulty levels, guaranteeing a representative mix. Additionally, we manually review each query before inclusion to ensure the properties of seed queries from Stage 1 are accurately preserved during the expansion process.
The resulting dataset follows a balanced difficulty distribution of 30% Easy, 40% Medium, and 30% Hard, facilitating fair cross-model comparisons while maintaining both challenge and diversity. It also upholds essential characteristics like implicit principles that can be concretely expressed through HTML and customized interaction.
4 Agentic Evaluation Methodology
As discussed in Section 2, assessing only static code or post-execution screenshots fails to verify interface behavior under real user interaction, nor to capture the implicit real-world principles required by the user’s query, which constitute two key challenges in generating high-quality MiniApps.
To address these challenges, MiniAppEval adopts an agentic evaluation framework with dynamic interaction enabled by browser automation (Playwright (Microsoft, 2026)). An LLM-powered agent actively interacts with the MiniApp and records the full interaction trajectory. Then based on this trajectory, MiniAppEval produces structured scores along three dimensions: Intention, Static, and Dynamic. Meanwhile, the evaluation framework is designed to minimize user cost. Users only need to provide an OpenAI-compatible chat API and can launch the entire evaluation with a single command (details in Appendix C.3; the cost analysis is provided in the Appendix C.6). For each query, the pipeline runs automatically, including code generation and scoring, which helps reduce the impact of extraneous factors unrelated to the model’s capabilities.
Overall, our methodology consists of two tightly coupled components: (i) a standardized code generation scaffold and (ii) an LLM-powered autonomous agentic evaluation framework.
4.1 Standardized Code Generation Scaffold
We provide an easy-to-use code generation scaffold. This part consists of two stages: Generation and Compilation.
Generation.
In the generation stage, the model receives a user query and generates a single, self-contained index.html file that integrates the document markup, embedded styling, and functional logic. Our evaluation uses the HTML format, while a standardized React option is also provided for users. The specific system prompt (generation prompt) templates are provided in Appendix E.1.
Compilation.
In the compilation stage, the generated source code is assembled and validated into a deployable artifact. All artifacts must be self-contained and runnable in a browser without external build tools, network access, or server-side dependencies. To ensure fair comparison, we run them in a standardized Chromium (Playwright) sandbox with fixed runtime conditions and strict isolation, evaluating each artifact independently.
4.2 Autonomous Agentic Evaluation Framework
Input.
The evaluation agent receives four inputs: (i) the original user query , (ii) the evaluation reference , (iii) the complete generated source code, and (iv) a live, interactable instance of the MiniApp running in the browser. Any natural-language explanations generated by the code model are retained as auxiliary context.
Evidence Collection.
MiniAppEval uses Playwright to simulate a human evaluator: it loads the generated MiniApp, observes its initial state, and autonomously interacts with it based on the user query . All interactions (clicking/typing) are executed via targeted JavaScript injected in the browser context for precise, deterministic control. The agent perceives rich signals (DOM, console logs, and source code; Appendix C.2.2) and selects actions (Appendix C.2.3) to probe functionality, guided by the query-specific evaluation reference to map requirements to verifiable checks and collect concrete evidence. The full process is recorded as a reproducible interaction trajectory (Appendix C.5).
Scoring.
Given the customized interactivity of MiniApps and their grounding in real-world principles, MiniAppEval combines static analysis with dynamic evidence to evaluate MiniApps along three dimensions: Intention, Static, and Dynamic. The evaluation reference , which encodes expected behaviors grounded in real-world principles, guides the agent’s inspection strategy but does not serve as a rigid oracle. Instead, the final judgment is based on whether the MiniApp functionally satisfies the user’s request. The output is a structured score across the three dimensions, each accompanied by a detailed rationale (highlighted in red at the top of Figure 4 (b)).
MiniAppEval departs from assertion-based or comparative benchmarks by directly evaluating whether a MiniApp satisfies open-ended user requirements, making it suitable for highly customized applications (the comparison shown in Figure 4).
Moreover, MiniAppEval addresses key limitations of human evaluation as shown on Figure 4 (b): (i) its static analysis precisely verifies implementation logic against real-world principles; (ii) Playwright’s programmatic control improves execution efficiency; and (iii) the LLM-powered evaluator leverages broad domain knowledge, often outperforming non-expert annotators on specialized tasks.
5 Experiments
| Model | Pass Rate (%) | Avg. (%) | Tokens | Time(s) | ||||||||
| Difficulty | Domain | |||||||||||
| Easy | Mid | Hard | Games | Science | Tools | Humanities | Viz. | Lifestyle | ||||
| Open-Source Large Language Models | ||||||||||||
1.59 | 0.55 | 0.00 | 0.00 | 0.57 | 0.00 | 0.00 | 2.04 | 3.70 | 0.66 | 3,470.68 | 22.16 |
6.43 | 2.35 | 0.00 | 0.93 | 0.60 | 4.00 | 4.88 | 7.27 | 10.34 | 2.88 | 4,068.27 | 49.55 |
6.06 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 9.43 | 11.11 | 1.83 | 2,324.83 | 25.04 |
14.17 | 5.03 | 0.00 | 3.77 | 3.11 | 4.08 | 4.88 | 17.65 | 18.52 | 6.19 | 3,435.97 | 46.76 |
17.60 | 4.07 | 1.44 | 5.66 | 4.27 | 6.98 | 7.32 | 16.98 | 10.34 | 7.09 | 7,110.65 | 58.94 |
36.30 | 15.06 | 4.41 | 12.50 | 10.49 | 20.00 | 17.07 | 35.19 | 48.39 | 18.31 | 8,936.88 | 55.58 |
| Closed-Source Large Language Models | ||||||||||||
6.32 | 0.87 | 0.00 | 0.00 | 0.00 | 3.03 | 0.00 | 13.51 | 3.57 | 2.32 | 3,727.55 | 132.67 |
28.68 | 8.33 | 2.22 | 13.46 | 6.02 | 10.87 | 11.63 | 23.53 | 36.36 | 12.48 | 5,109.82 | 37.98 |
29.66 | 12.12 | 2.19 | 8.41 | 6.58 | 20.00 | 17.50 | 32.65 | 25.93 | 13.77 | 9,010.00 | 75.62 |
31.46 | 15.62 | 7.08 | 16.25 | 12.50 | 23.33 | 20.00 | 27.27 | 19.23 | 17.12 | 8,881.57 | 118.32 |
32.76 | 16.89 | 4.10 | 14.95 | 10.60 | 17.95 | 18.18 | 30.61 | 41.38 | 17.62 | 6,563.28 | 50.56 |
61.98 | 20.83 | 1.71 | 26.74 | 19.11 | 13.64 | 28.57 | 52.00 | 55.56 | 27.52 | 5,815.14 | 80.80 |
68.22 | 14.86 | 1.79 | 16.13 | 22.30 | 29.27 | 23.81 | 47.73 | 44.83 | 26.36 | 8,586.84 | 91.43 |
59.09 | 41.18 | 22.33 | 37.18 | 34.59 | 47.50 | 35.71 | 57.45 | 56.52 | 41.14 | 13,152.75 | 166.66 |
74.71 | 21.37 | 3.49 | 24.14 | 18.10 | 33.33 | 45.83 | 57.78 | 64.71 | 32.00 | 11,256.15 | 154.09 |
69.77 | 43.08 | 18.64 | 40.32 | 50.38 | 50.17 | 45.45 | 75.00 | 82.35 | 45.46 | 10,793.68 | 169.60 |
| Average | 34.05 | 13.89 | 4.34 | 14.71 | 11.64 | 18.07 | 17.55 | 31.63 | 33.30 | 17.05 | – | – |
5.1 Settings
All evaluations are conducted in a sandbox with deterministic seeds and fixed rendering settings. Artifacts are rendered via Playwright (headless Chromium) at multiple resolutions, including 1280720, to test adaptive designs. Models receive identical prompts (listed in Appendix E.1) and follow a unified decoding protocol: we use officially recommended decoding parameters when available; otherwise, we apply our defaults (detailed in Appendix C.1.1). Overlong inputs are truncated, and each run is capped at 15 minutes.
Baseline models are selected to ensure breadth, currency, and reproducibility, considering: (1) multiple model families (Claude (Anthropic, 2025a, b), Gemini (Google DeepMind, 2025a, b), GLM (Zeng et al., 2025), GPT (OpenAI, 2025b, a), Grok (xAI, 2025), Hunyuan (Team et al., 2025b), Kimi (Team et al., 2025a), Mimo (Xiao et al., 2026), MiniMax (Chen et al., 2025), and Qwen3 (Yang et al., 2025)); (2) a range of scales (from lightweight to flagship); and (3) relatively recent and representative versions within each family. For evluation, we select Gemini-3-pro as the evaluation-model driving agent due to its strong agreement with human judgments.
5.2 Main Results and Analysis
Our framework supports custom thresholds. In our experiments, we adopt a threshold of 0.8: a MiniApp is considered successful if its minimum score across the three dimensions (Intention, Static, Dynamic) exceeds this value, i.e., . GPT-5.2 achieved the highest performance with an average pass rate of 45.46%, while the overall mean across all models was 17.05%. These results underscore the challenges current models face in generating successful MiniApps. The details are shown in Figure 5 and Table 1.
Open-Source vs. Closed-Source Performance Analysis.
Our experiments show a clear gap between open- and closed-source models, with closed-source systems consistently performing better across all difficulty levels. In contrast, benchmarks such as ArtifactsBench (Zhang et al., 2025) and WebDevJudge (Li et al., 2025a) report much smaller gaps, suggesting potential saturation or overfitting; our benchmark better avoids this issue and thus provides a more discriminative evaluation.
Difficulty-Level Performance Analysis.
The difficulty-wise performance analysis validates the rationale behind our task difficulty gradient segmentation, showing that models with different performance levels can find their respective niches when tackling tasks of varying complexity. As shown on Table 1, the accuracy of all models decreases with increasing difficulty. Furthermore, smaller open-source models (Qwen3-32B) can handle certain tasks effectively, whereas more advanced models often struggle with more complex challenges.
Domain-wise Performance Analysis.
As shown on Table 1, the performance varies significantly across different classes. The pass rates for the Visualization and Lifestyle categories are notably higher, exceeding 30%, with GPT-5.2 performing particularly well. This suggests that current models excel in tasks with a clear, singular objective, such as visualizations, and in tasks that just require the application of commonsense. However, for more complex categories that involve comprehensive tasks, domain-specific knowledge, and intricate engineering details, the models still exhibit some limitations.
Model-Scale and Positioning Analysis.
Across both the Qwen and GLM families, we observe a consistent trend where increasing model scale generally leads to superior performance, validating the impact of scaling laws on complex tasks. Within the Qwen3 series, Qwen3-235B-A22B achieves a 2.88% pass rate, significantly outperforming the smaller Qwen3-32B (0.66%). This scaling trajectory is even more pronounced in the GLM series: the lightweight GLM-4.5-Air achieves a 7.09% pass rate, while the flagship GLM-4.7 reaches a substantial 18.31%, illustrating the massive performance gains derived from increased model capacity and architectural refinement.
Performance vs. Inference Cost Analysis.
There is a strong positive correlation between performance and token consumption (0.8433), and a moderate correlation with time (0.7387), as illustrated in Figure 6, suggesting that more tokens and time generally improve performance. The correlation is measured by the Pearson correlation coefficient (Pearson, 1895). Outliers include GPT-5.2 and Gemini-3-Pro-Preview, which consume fewer tokens than models with similar performance. Hunyuan-Turbos-Latest and MiniMax-M2.1 have notably higher processing times for similar performance.
5.3 Ablation study
To evaluate the impact of different components on the performance of MiniAppEval, we conducted an ablation study on a set of 183 manually labeled ground truth (GT) samples, as shown in Table 2.
| Exp. | Acc. | Prec. | Rec. | F1 |
| MiniAppEval | 89.62 | 83.87 | 85.25 | 84.55 |
| w/o Code | 70.66
18.96 |
32.73
51.14 |
60.00
25.25 |
42.35
42.20 |
| w/o Agent | 66.48
23.14 |
12.90
70.97 |
53.33
31.92 |
20.78
63.77 |
| w/o Eval Ref | 60.12
29.50 |
89.47
5.60 |
46.36
38.89 |
61.08
23.47 |
The full MiniAppEval system (comprising Eval-Ref, Code, and Playwright) achieves the highest accuracy among all variants, demonstrating the overall effectiveness of the proposed evaluation framework. Removing the Eval-Ref leads to a substantial drop in recall, indicating that the Eval-Ref plays a critical role in guiding MiniAppEval to attend to the correct aspects of a query and to accurately localize potential failure cases. w/o Code results in a sharp degradation in precision, as the judge can no longer verify implementation details (e.g., detect violations of implicit real-world principles). w/o Agent yields the lowest precision overall, highlighting that many interaction-dependent behaviors can only be revealed through active exploration, which are inaccessible to static inspection alone.
5.4 Double Blind Judge
During evaluation, we observed that for graphical queries (e.g., in the Visualization class), the agent judge could be overly lenient due to confirmation bias (Nickerson, 1998). To mitigate this, we introduce a double-blind evaluation procedure (detailed in Appendix D): the judge first evaluates the output without seeing the query, and then checks it against the user requirements for the final decision. We apply this protocol to 55 graphical queries. As shown in Table 3, it improves accuracy and better identifies negative samples, supporting our hypothesis and offering a more reliable setup for purely visual tasks.
| Model | Method | T/T | T/F | F/T | F/F | Acc. |
| Gemini-3-Pro | MiniAppEval | 15 | 2 | 8 | 30 | 81.82 |
| -Pro-Preview | Double-Blind | 11 | 6 | 2 | 36 | 85.45
3.63 |
| GPT-5.2 | MiniAppEval | 16 | 3 | 8 | 28 | 80.00 |
| Double-Blind | 12 | 7 | 2 | 34 | 83.63
3.63 |
| Claude- | MiniAppEval | 17 | 3 | 9 | 26 | 78.18 |
| Opus-4.5 | Double-Blind | 11 | 9 | 0 | 35 | 83.63
5.45 |
5.5 Validation of Evaluation Effectiveness
To validate the effectiveness and reliability of MiniAppEval, we conducted a human agreement study with four experts on 183 items from each of three representative models spanning different performance tiers: low- (GLM-4.7), mid- (Gemini-3-pro-preview), and high-performing (GPT-5.2) (549 outputs total); each output was annotated by all four experts (2,196 annotations).
We first assessed inter-rater reliability using Fleiss’ Kappa (Fleiss, 1971), obtaining . Using the aggregated expert labels as reference, we then computed Cohen’s Kappa (Cohen, 1960) between MiniAppEval and humans across the three models to cover different quality regimes. As shown in Table 4, MiniAppEval achieves strong agreement with humans, with ranging from 0.81 to 0.89.
| Model | TP | FP | FN | TN | acc | Cohen’s |
| Gemini-3-pro-preview | 83 | 8 | 9 | 83 | 0.9071 | 0.8142 |
| GLM-4.7 | 87 | 5 | 5 | 86 | 0.9454 | 0.8907 |
| GPT-5.2 | 85 | 7 | 7 | 84 | 0.9235 | 0.8470 |
6 Conclusion
In conclusion, we introduce MiniAppBench, the first benchmark for evaluating principle-driven interactive application generation, addressing key gaps left by prior benchmarks. We further propose MiniAppEval, an agentic, browser-based evaluation framework that enables comprehensive and automated assessment of MiniApps. Our experiments show that current LLMs still struggle to generate high-quality MiniApps, while MiniAppEval aligns closely with human judgments, providing a reliable method for future research.
Appendix A Benchmark Comparison
To position MiniAppBench among existing evaluations, we compare representative benchmarks from three research lines: code generation, real-world reasoning, and web development. Table 5 summarizes their data scale, task type, real-user sourcing, and the degree to which they require principle-grounded interactive behavior. Notably, MiniAppBench is the first benchmark that integrates real-user queries, high task diversity/complexity, and explicit real-world principle requirements into a single unified evaluation setting.
| Benchmark | #Data | Task | Real-User | Div. | Comp. | RW-Prin. |
| MBPP | 500 | Algorithmic Problem Solving | ✗ | Low | High | Low |
| HumanEval | 164 | Algorithmic Problem Solving | ✗ | Low | High | Low |
| SWE-Bench | 2,294 | Repository-level Bug Fixing | ✗ | High | High | Low |
| MLE-Bench | 75 | Repository-level Software Engineering | ✓ | High | High | Low |
| PIQA | 2,000 | Physical Reasoning | ✗ | Low | Low | High |
| GSM8K | 1,000 | Mathematical Reasoning | ✗ | Low | Low | High |
| AlfBench | 3,553 | Embodied Reasoning | ✗ | Low | Low | High |
| Voyager | N/A | Embodied Reasoning | ✗ | Low | Low | High |
| Pix2Code | 5,250 | Web Interface Cloning | ✗ | Low | Low | Low |
| Web2Code | 1,198 | Web Interface Cloning | ✗ | Low | Low | Low |
| FullFront | 50 | Web Interface Cloning | ✗ | Low | High | Low |
| WebGenBench | 101 | Multi-file Web Dev | ✗ | Mid | High | Low |
| A11YN | 300 | Web Accessibility | ✓ | High | Low | Low |
| WebBench | 50 | Multi-step Iterative Dev | ✗ | High | High | Low |
| FronTalk | 100 | Multi-step Iterative Dev | ✗ | Low | High | Low |
| ArtifactsBench | 1,825 | Interactive Visual Artifacts Dev | ✗ | High | Mid | Mid |
| WebDevArena | N/A | Web Preference (A/B) | ✓ | High | – | – |
| WebDevJudge | 654 | Web Preference (A/B) | ✓ | Mid | High | Mid |
| MiniAppBench | 500 | Customized MiniApps Dev | ✓ | High | High | High |
Appendix B Data Construction
B.1 Domain Classification
The two-level taxonomy of queries is carried out in two stages. First, we utilize large models to categorize the queries based on real-world user data, generating an initial classification. Subsequently, human experts review and refine this categorization, abstracting a more logical and coherent classification scheme. The final classification consists of six coarse-grained domain: Science, Games, Tools, Humanities, Lifestyle, and Visualization. Each coarse-grained domain is further subdivided, with specific subclass outlined in the Table 6.
To ensure a comprehensive evaluation of model capabilities, we also considered the proportional distribution of categories when constructing the dataset. Using the real-world query distribution as a baseline, we made necessary adjustments to maintain a reasonable balance across categories (for instance, due to the higher demand for game-related queries in online data, we reduced the proportion of game-related queries but still kept it significant).
| Domain | Subclass | Count | Ratio (%) |
| Science | Chemical | 46 | 9.20 |
| Biological Systems | 44 | 8.80 | |
| Physics | 37 | 7.40 | |
| Virtual Laboratory | 35 | 7.00 | |
| Geometry | 25 | 5.00 | |
| Total (Science) | 187 | 37.40 | |
| Games | Logic | 28 | 5.60 |
| Projectile | 25 | 5.00 | |
| Reflex | 16 | 3.20 | |
| Edutainment | 16 | 3.20 | |
| Systemic Simulation | 15 | 3.00 | |
| Casual | 11 | 2.20 | |
| Card | 10 | 2.00 | |
| Total (Games) | 121 | 24.20 | |
| Tools | Schedule | 21 | 4.20 |
| Creative Tools | 18 | 3.60 | |
| Computational Tools | 15 | 3.00 | |
| Data Lookup | 3 | 0.60 | |
| Total (Tools) | 57 | 11.40 | |
| Humanities | Skill Acquisition | 26 | 5.20 |
| Concept Deconstruction | 13 | 2.60 | |
| Culture | 8 | 1.60 | |
| Total (Humanities) | 47 | 9.40 | |
| Lifestyle | Health | 14 | 2.80 |
| Toys | 10 | 2.00 | |
| Roleplay | 8 | 1.60 | |
| Total (Lifestyle) | 32 | 6.40 | |
| Visualization | SVG | 25 | 5.00 |
| Statical | 23 | 4.60 | |
| Art | 8 | 1.60 | |
| Total (Visualization) | 64 | 11.20 | |
| Grand Total | 500 | 100.00 | |
B.2 Screening Guidelines
B.2.1 Customized Interaction
To ensure that MiniAppBench targets customized interaction rather than conventional template-driven web development, we screen candidate queries by checking whether the requested behavior requires synthesizing query-specific interaction logic that cannot be reduced to standard CRUD workflows (e.g., form submission database update list rendering).
Concretely, a query is labeled as requiring customized interaction if it satisfies at least one of the following criteria:
Multi-step state transitions. The task requires maintaining and updating non-trivial internal states across multiple user actions (e.g., “simulate one week of choices”, “step-by-step experiment”, “undo/redo”, “scenario branching”), beyond add/edit/delete of records.
Custom interaction operators. The task involves interaction primitives that are not typical CRUD UI patterns, such as dragging, drawing, manipulating sliders to control a simulation, playing a game, interactive diagram exploration, timeline scrubbing, or parameter sweeping.
Dynamic rules grounded in the query. The runtime behavior must obey explicit or implicit rules that are unique to the query, such as physical laws (gravity, conservation), temporal constraints (a week has seven days), geometric constraints, scoring rules in a game, or procedural generation rules.
Open-ended user exploration. The user is expected to explore a concept by interacting with the interface (e.g., “interactive visualization to understand …”, “what-if analysis”), where the value arises from the interaction trajectory rather than static content display.
Non-trivial edge-case handling. The query implies boundary conditions that affect interaction logic (e.g., invalid parameter ranges, impossible states, constraint violations) and thus requires tailored runtime checks beyond form validation.
We exclude queries that can be adequately solved by: (i) static information presentation (e.g., “show me an introduction to …”), or (ii) standard CRUD-style applications (e.g., “create a webpage to add/edit/delete notes”), where the interaction can be implemented with a generic form-list template and does not require query-specific dynamics.
During screening, each candidate query is independently reviewed by two annotators following the above criteria. Disagreements are resolved through discussion, and borderline cases are retained only if the interaction logic is clearly driven by query-specific rules rather than templated CRUD patterns.
B.2.2 Real-world Principle
In addition to customized interaction, we require each query to involve at least one real-world principle that constrains the MiniApp’s behavior. Here, a principle refers to an implicit or explicit rule about how the world should work (e.g., physical laws, temporal constraints, domain conventions, or commonsense invariants) that must be operationalized in an executable artifact.
Principle taxonomy. Our principle categorization follows the European Research Area (ERA), covering four broad areas: Life Sciences, Physical Sciences and Engineering, Social Sciences and Humanities, and Health and Medicine. Each query is annotated with the area(s) of principle it primarily relies on (e.g., conservation laws in a physics simulation; biological processes in a cell-cycle demo; historical timelines and causal narratives in humanities; dosage/health constraints in medicine).
HTML-expressibility requirement. Crucially, we only retain queries whose underlying principles can be faithfully expressed and verified through a browser-executable interface. Our screening assumes the following executable-web decomposition: HTML represents world states and structural relationships, CSS determines perceptual salience, and JavaScript encodes causal dependencies, temporal evolution, and interaction logic—together forming an executable world model. Therefore, a query passes the principle screening only if the principle can be mapped to at least one of the following HTML-expressible forms:
State representation: the relevant entities, attributes, and constraints can be represented as DOM elements and state variables (e.g., positions, counts, schedules, scores).
Rule execution: the principle can be implemented as deterministic or stochastic update rules in JavaScript that govern state transitions over time and user interactions (e.g., numerical integration for motion, discrete event simulation, rule-based scoring).
Perceptual grounding: the principle’s outcomes can be rendered and inspected via visual encodings or UI feedback (e.g., trajectories, charts, alerts, invariants displayed as diagnostics).
We exclude queries whose required principles are not meaningfully capturable in an offline, self-contained browser setting, such as tasks requiring external sensors, proprietary databases, real-time web access, or unverifiable claims that cannot be grounded in executable state-transition logic. This ensures that every retained query admits a MiniApp implementation where principle adherence is both implementable and testable within HTML/CSS/JavaScript.
B.3 Data Format
The evaluation dataset is stored as a JSON array. Each element corresponds to one MiniApp specification and its LLM-generated evaluation reference. Each record contains six fields: index, class, subclass, query, level, and eval-reference (a JSON-serialized string).
Fields. index is a unique identifier (1-based) within the file. class and subclass denote the coarse- and fine-grained categories. query is the natural-language specification used for generation. level is the difficulty tag (Easy/Mid/Hard). eval-reference encodes the evaluation reference in three dimensions (intention, static, dynamic) and is parsed by the evaluator when needed.
Appendix C MiniAppEval
C.1 Settings
C.1.1 Models’ Decoding Protocol
We follow each model’s official API documentation or default demo settings when available. For models without explicit recommendations, we adopt commonly used default values to ensure fair comparison. The specific settings shown on Table 7
| Model | Temperature | Top- | Max tokens |
| GPT-5.2 | 1.0 | 1.0 | 128,000 |
| GPT-5.1 | 1.0 | 1.0 | 400,000 |
| Claude-Opus-4.5 | 1.0 | 1.0 | 200,000 |
| Claude-Sonnet-4.5 | 1.0 | 1.0 | 200,000 |
| Gemini-3-Pro-Preview | 0.8 | 0.95 | 65,536 |
| Gemini-3-Flash | 0.8 | 0.95 | 65,536 |
| GLM-4.7 | 1.0 | 0.95 | 131,072 |
| GLM-4.5-Air | 1.0 | 0.95 | 96,000 |
| MiniMax-M2.1 | 1.0 | 1.0 | 204,800 |
| Grok-4.1-Fast-Reasoning | 1.0 | 1.0 | 30,000 |
| Mimo-V2-Flash | 1.0 | 1.0 | 32,768 |
| Kimi-K2-Instruct | 1.0 | 1.0 | 256,000 |
| Qwen3-235B-A22B | 1.0 | 1.0 | 38,912 |
| Qwen3-Coder-480B-A35B-Instruct | 1.0 | 1.0 | 65,536 |
| Qwen3-32B | 1.0 | 1.0 | 32,768 |
| Hunyuan-Turbos-Latest | 1.0 | 1.0 | 256,000 |
C.1.2 Two MiniApps Generation Formats
To more comprehensively evaluate model capabilities while reducing interference from output formatting, our evaluation supports two generation modes: (1) a single-file HTML mode and (2) a React framework mode. For both modes, the pipeline automatically extracts the generated code, builds a runnable project, launches it in a sandboxed environment, and then completes the evaluation. The recommended file structure for the React mode is shown below. Prompts for both generation formats are provided in E.1.
C.1.3 Positive/Negative Labeling
We convert the three-dimensional evaluation scores into a binary label for downstream analysis. A MiniApp is marked as positive (successful) if all three dimension scores exceed a predefined threshold, i.e.,
(we use in the main setting). Otherwise, it is labeled as negative (failed). This conservative rule ensures that a sample is counted as successful only when it simultaneously satisfies the user intention, static correctness, and dynamic interaction requirements.
C.2 Environment Setup
Our evaluation framework conducts agent assessments in web-based environments, enabling comprehensive evaluation of GUI agents through automated browser interaction and code analysis. The evaluation system is implemented through a standardized evaluation script that provides a consistent interface for assessing agent-generated web applications. In the following sections, we detail the environment design for web-based agent evaluation.
C.2.1 Environment Infrastructure
We design an interactive web-based evaluation environment using browser automation technology. The environment leverages Playwright as the browser automation platform through the Model Context Protocol (MCP) server interface, enabling high compatibility with real-world web applications while maintaining full control over the execution environment. This setup allows us to simulate user interactions such as mouse clicks, keyboard input, and form submissions, which are essential for evaluating GUI agents’ capabilities. The browser automation framework supports real-time observation and logging of DOM states, facilitating fine-grained analysis and reproducibility of agent behavior. All evaluation episodes are initialized from a clean browser state to ensure consistent starting conditions for each evaluation episode. The evaluation system supports two complementary modes: standard mode, where agents interact with live web applications via URLs with full browser automation capabilities, and code-only mode, where evaluation is performed solely based on HTML and JavaScript code analysis without browser access. This dual-mode design enables flexible evaluation strategies, allowing assessment of both runtime behavior and static code quality.
C.2.2 Observation Space
In our evaluation framework, the observation space is designed to ensure comprehensive evaluation of web-based GUI agents by capturing both structural and semantic aspects of web pages. It comprises two complementary modalities: DOM structure snapshots and source code access. The DOM snapshot is obtained through the Playwright MCP server’s browser_evaluate interface, which provides a complete representation of the page’s hierarchical structure, including all HTML elements, their attributes, text content, and accessibility information. This structural information enables agents to understand the page layout, identify interactive elements, and navigate the interface effectively. Additionally, when available, agents can access the HTML and JavaScript source code directly, which provides insights into the implementation details, event handlers, and application logic. This dual-modality approach reflects the varying capabilities of different agent architectures. For example, agents that have been specifically trained on web environments often possess strong grounding abilities and can rely on DOM snapshots alone. In contrast, general-purpose language models typically benefit significantly from the semantic and structural information provided by both DOM structure and source code. By supporting both modalities, our framework enables fair and informative evaluation across a wide range of agents, ensuring robust assessment under diverse web application contexts and UI layouts. Notably, the framework explicitly prohibits the use of visual screenshots or rendering-based analysis, focusing exclusively on structural and semantic information to ensure objective and reproducible evaluation.
C.2.3 Action Space
In our evaluation framework, the action space consists of core types of user interactions that an agent can perform to interact with web applications. These actions, summarized in Table 8, enable the agent to effectively interact with graphical user interfaces across a wide range of web applications.
| Action | Description |
| browser_click | Simulates mouse clicks on UI control elements. Supports configurable mouse buttons (left, right, middle) and both single and double clicks. Commonly used for selecting items, activating controls, or triggering events. |
| browser_type | Simulates keyboard input for entering text, pressing keys, or invoking shortcuts (e.g., Ctrl+C, Enter). Enables fine-grained control over application behavior and supports both functional input and text entry. |
| browser_fill_form | Fills form fields with specified values, supporting various input types including text inputs, checkboxes, radio buttons, and dropdown selections. Allows batch form filling for efficient interaction with complex forms. |
| browser_evaluate | Executes JavaScript code to query DOM state or perform complex operations. Enables agents to extract information, manipulate page elements, or verify application state programmatically. Particularly useful for analyzing CSS styles, color schemes, and dynamic content. |
| browser_wait_for | Waits for specific conditions such as element appearance, text changes, or custom JavaScript predicates. Essential for handling asynchronous operations and ensuring elements are ready before interaction. |
This comprehensive action space allows agents to perform complex multi-step interactions, test dynamic behaviors, and verify application functionality across diverse web application scenarios. The combination of basic interaction actions (browser_click, browser_type, browser_fill_form) with advanced programmatic capabilities (browser_evaluate, browser_wait_for) enables thorough evaluation of both static UI elements and dynamic interactive behaviors. The framework emphasizes that all interactions must be verified through actual DOM state changes rather than assumptions, ensuring that evaluation results reflect genuine application capabilities rather than inferred behavior.
C.3 The Pipeline of Agentic Evaluation
We design a one-click evaluation pipeline. Given only an OpenAI-compatible API endpoint, the system automatically runs the entire workflow, including loading queries, generating MiniApps, and evaluating the generated artifacts, while recording detailed logs. It supports multiple modes: generation can be performed in either HTML mode or React mode; evaluation includes, but is not limited to, MiniAppEval, evaluation without code access, and evaluation without evaluation references. The pipeline also supports batched execution, substantially reducing evaluation overhead. Moreover, by standardizing both the generation scaffold and the evaluation environment, it minimizes external confounding factors and improves the fairness of experimental results. The pseudo-code of the workflow is shown below.
C.4 Results Format
For each generated MiniApp, the evaluator produces a structured JSON result with three dimensions: intention, static, and dynamic. Each dimension contains (i) a scalar score in and (ii) a short natural-language reason explaining the judgment. The overall pass/fail decision in our experiments is derived from these three scores (see C.1.3 for the thresholding rule), while the reason fields are retained for error analysis and qualitative inspection (Example in below).
C.5 Evaluation Trajectory
An evaluation trajectory records the step-by-step execution of the agent during MiniAppEval, including the conversation context, model outputs, tool calls, and token/time usage. Trajectories are stored as JSONL files, where each line corresponds to one evaluation step.
[…]
C.6 Time, Token Consumption, and Step Analysis
We analyze the trajectory logs collected by MiniAppEval over 44,981 valid runs. Figure 7 provides a compact, multi-view visualization of the relationships among step count, token consumption, and latency. Overall, we observe three consistent patterns: (i) token usage increases mildly with step progression, largely due to accumulated prompt context; (ii) per-step time intervals exhibit substantial variance and a long-tailed distribution; and (iii) prompt tokens dominate the overall token budget, while completion tokens account for only a small fraction. These findings suggest that evaluation cost is primarily driven by interaction length and context growth, and motivate future optimizations in context management and evaluation efficiency.
Appendix D Double Blind Evaluation
D.1 Experimental Design
The double-blind evaluation method addresses confirmation bias (Nickerson, 1998) by separating objective observation from subjective judgment through a two-stage process. This approach is particularly effective for graphical queries in visualization tasks, where evaluators may exhibit leniency due to cognitive bias when directly comparing implementations against queries.
The evaluation workflow consists of two sequential stages:
Stage 1 (Blind Description): An agent is provided with only the webpage code (excluding descriptive text) and a URL, without access to the user query. The agent generates a structured, objective description of the page’s visual and interactive elements.
Stage 2 (Consistency Scoring): A separate evaluation model receives the Stage 1 description output along with the original query and optional evaluation reference. Based solely on the description rather than direct page access, the model generates a consistency score and detailed analysis.
D.2 Experimental Results
We conducted experiments comparing double-blind evaluation with the standard evaluation method on a test set of 55 graphical queries. Three models (Gemini-3-Pro, GPT-5.2, and Claude-Opus-4.5) were used to generate evaluation targets, and both evaluation methods were applied to each set of results.
The experimental results demonstrate that double-blind evaluation achieves higher accuracy and effectively mitigates confirmation bias. Specifically, double-blind evaluation achieved an average accuracy of 84.24%, compared to 80% for the standard method. More importantly, for manually labeled negative samples, double-blind evaluation showed significantly higher accuracy (96.33% vs. 77.06%), indicating greater sensitivity to negative cases and effective elimination of cognitive bias introduced by query context.
However, for positive samples, double-blind evaluation showed lower accuracy (60.7% vs. 87.27%), suggesting a more stringent evaluation standard. This stricter approach further validates that standard evaluation methods are constrained by confirmation bias, where evaluators may adjust their expectations to match observed implementations rather than maintaining objective assessment criteria.
The two-stage design ensures that Stage 2 evaluators cannot access the original webpage and must reason entirely from the Stage 1 description. This constraint forces evaluators to work with factual observations rather than making assumptions, thereby reducing the tendency to retroactively align expectations with implementations. The structured format of stage1_description ensures consistency across different observers, while stage2_evaluation provides both quantitative scores and qualitative reasoning for interpretability.
Appendix E Prompts
In this section, we present the prompts used for generating MiniApps, evaluating and building evaluation reference.
E.1 Prompts for Generating MiniApps
E.2 Prompts for MiniAppEval
E.3 Prompts for Double Blind Judge Evaluation
E.4 Prompts for Building Evaluation Reference
Since the evaluation reference is a core component of our benchmark—directly defining the scoring criteria and alignment—releasing the exact prompts used to construct it could encourage “teaching to the test” and lead to overfitting. To preserve fairness, robustness, and the validity of our evaluation results, we therefore choose not to disclose the prompts in this subsection in the current release. We will consider sharing additional details in a safer form in future versions without compromising the integrity of the benchmark.